# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v1

@@ Introduction
Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al. 2020, 2022). Details are shown in Appendix A3. Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively. The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$.
@@
Algorithm 1 SAMBO-RL
Input: Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, shifts-aware reward hyperparameters $\alpha$ and $\beta$.
3:  for $i=1,2,...,N_{iter}$ do
5:     Generate synthetic rollouts by model $m_{\theta}$. Add transition data in these rollouts to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
6:     Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (7).
7:     Compute shifts-aware reward according to Eq. (8).
8:     Run SAC (Haarnoja et al. 2018) with shifts-aware reward to update policy $\pi$.
9:  end for
Experiments
@@
We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance. Our implementation is based on the OfflineRL-Kit library***https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of our algorithm are consistent with the settings in this library.
Effectiveness of Shifts-aware Reward
@@
We conduct an evaluation of the effectiveness of the Shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid environment illustrated in Fig. 1 (a).
Model Bias Adjustment.
Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective. To evaluate the effectiveness of model bias adjustment, we examine its amendment to the training process. As illustrated in Fig. 2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM), the shifts-aware reward with the overestimating model (SAR with OM), the vanilla reward with the underestimating model (UM), and the shifts-aware reward with the underestimating model (SAR with UM). We then measure the expected return of the learned policy after each update during training in the actual environment. The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy. In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, thereby sustaining optimal returns. Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving near highest return.
Policy Shift Modification.
Policy shift implicitly restricts policy updates. To assess the effectiveness of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig. 1 (a). We test four settings: using either the vanilla reward or the shifts-aware reward, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy. We measure the KL divergence between the learned policy and the behavior policy after each update. As illustrated in Fig. 2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence. In contrast, the algorithm employing the shifts-aware reward effectively motivates policy to be aggressive and optimistic and overcomes this restriction, displaying a higher and more stable divergence and achieving faster policy convergence.
Results on the D4RL Benchmark
Datasets.
In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al. 2020) on the MuJoCo simulator (Todorov, Erez, and Tassa 2012). The evaluation encompasses 12 datasets spanning three environments (HalfCheetah, Hopper, Walker2d) and four dataset types (random, medium, medium-replay, medium-expert). The “v2” versions of datasets are employed for assessment.
Baselines.
We compare SAMBO against multiple baselines, including both model-free and model-based offline RL algorithms. For model-free approaches, BC (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011) learns behavior policy, CQL (Kumar et al. 2020) penalizes the Q-function according to policy change, TD3+BC (Fujimoto and Gu 2021) adds a BC term to regularize the policy, and EDAC (An et al. 2021) quantifies the uncertainty of the Q-value using an ensemble of Q-function networks. For model-based approaches, MOPO (Yu et al. 2020) penalizes rewards via the predicted variance of models. COMBO (Yu et al. 2021) implements CQL within a model-based framework. TT (Janner, Li, and Levine 2021) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter, Lacerda, and Hawes 2022) adversarially trains the policy and the dynamics model within a robust framework. MOBILE (Sun et al. 2023) penalizes the value learning of model data based on the uncertainty of the Bellman Q-function, derived from an ensemble of models.
Comparison Results.
Results are presented in Table 1. Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al. 2020). SAMBO achieves the highest scores on two random datasets and demonstrates strong performance across both medium and medium-replay datasets. Additionally, SAMBO shows competitive results on medium-expert datasets. This performance aligns with our theoretical analysis. By addressing both model bias and policy shift, SAMBO effectively balances the exploitation of offline data with the need to overcome the limitations of the behavioral policy. This approach enables the algorithm to be more aggressive and optimistic in its search for the optimal policy. In scenarios involving low-quality datasets, the aggressive exploration of SAMBO is more likely to yield benefits, potentially leading to the discovery of a superior policy. Conversely, a conservative approach may impede the learning of optimal policy in such contexts. For high-quality datasets, the advantages of aggressive exploration tend to diminish. Fortunately, in these scenarios, methods that emphasize estimating Q-value uncertainty perform effectively. This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
@@
walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 82.8$\pm$0.2
halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 94.8$\pm$2.8
hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 106.5$\pm$5.6
walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 80.2$\pm$6.8
Ablation Studies
@@
(a) Ablation study for SAMBO
(b) Hyperparameter tuning for $\alpha$
(c) Hyperparameter tuning for $\beta$
Figure 3: Illustration of hyperparameter experiments in the HalfCheetah task with a random dataset. (a) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward. (b) A large model bias adjustment coefficient $\alpha$ increases conservativeness. (c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
Ablation.
@@
Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification. To elucidate the importance of these components, we perform ablation studies on the Halfcheetah task. We assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent. The ablation results are presented in Figure. 3 (a).
@@
We find that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates. This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration. However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance. In contrast, SAMBO incorporates both model bias adjustment and policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance. This underscores the effectiveness of model bias adjustment.
The ablation experiment wo/ps incorporates model bias adjustment but does not address policy shift. As a result, while wo/ps mitigates value misestimation caused by model bias, the lack of policy shift modification causes the algorithm to remain implicitly constrained by the behavioral policy. Therefore, wo/ps demonstrates reduced effectiveness in discovering a more optimal policy. In contrast, SAMBO integrates policy shift modification during conservative updates, which enables more aggressive exploration and results in superior performance compared to wo/ps. This highlights the effectiveness of incorporating policy shift modification.
Hyperparameter Tuning.
SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
The model bias adjustment coefficient $\alpha$ regulates the extent of adjustment applied to model-generated data. A larger $\alpha$ imposes a significant penalty on the model, leading to excessively conservative utilization of the model. Consequently, even in regions where the model performs well, exploration may be restricted due to this excessive penalty, potentially resulting in suboptimal performance. Conversely, a smaller $\alpha$ can cause the algorithm to be overly optimistic about exploring risky regions, thereby impairing performance. To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task. The results, shown in Fig. 3 (b), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance. Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration. This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
@@
The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy. A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy. However, this also introduces a heightened risk of taking unsafe actions. As illustrated in Fig. 3 (c), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness. A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability. Therefore, $\beta$ also needs to be set appropriately to balance exploration and stability.
Conclusion and Discussion
@@
In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift. Model bias results in distorted objectives, while policy shift obstructs the convergence of policies. We introduce a shifts-aware policy optimization framework to uniformly address these challenges and apply it within the model-based offline setting. Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms. Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training. In future work, We aim to explore alternative methods that directly address distribution shift using datasets, eliminating the need for classifier estimation. Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems. Our novel analysis and results related to policy shift in this context open new avenues for better exploiting offline datasets and overcoming the restrictions imposed by behavior policies.
Appendix A A. Theoretical Analysis
A1. The Supplement Analysis of Distribution Shift

### v2

@@ 4.3 Shifts-aware Model-based Offline RL
Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al., 2020, 2022). Details are shown in Appendix A.3. Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively. The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$.
@@
5 Experiments
@@
We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance. Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of our algorithm are consistent with the settings in this library.
5.1 Effectiveness of Shifts-aware Reward
@@
We conduct an evaluation of the effectiveness of the shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid world illustrated in Fig. 1 (a).
@@
walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 92.3$\pm$ 1.5
halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 97.8 $\pm$0.3
hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 111.0 $\pm$1.2
walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 106.1 $\pm$ 2.9
Model Bias Adjustment. Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective. To evaluate the effects of model bias adjustment on the training process, we conduct experiments. As illustrated in Fig. 2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM) or underestimating model (UM), the shifts-aware reward with the overestimating model (SAR with OM) underestimating model (SAR with UM). We then evaluate the expected return of the learned policy after each training update in the actual environment. The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy. In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, sustaining optimal returns. Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving a return near the theoretical optimum.
(a) MOBILE
(b) SAMBO
Figure 3: The $(s,a,s^{\prime})$ distributions of offline RL algorithms result in the Hopper-M task. Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
Policy Shift Modification. Policy shift implicitly restricts policy updates. To assess the efficacy of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig. 1 (a). We test four settings: the vanilla reward or SAR, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy. We measure the KL divergence between the learned policy and the behavior policy after each update. As illustrated in Fig. 2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence. In contrast, the algorithm employing the SAR effectively motivates policy to be optimistic. It overcomes the restriction, displaying a higher and more stable divergence and achieving faster convergence.
Effectiveness in Complex Tasks. We also conducted experiments on complex tasks. Fig. 3 shows a visual comparison in the NeoRL task (details in Sec. 5.2.2). SAMBO effectively constrains model-generated data (i.e., Model Data) to align closely with the Offline Data, whereas the previous state-of-the-art method, MOBILE, imposes weaker constraints. Furthermore, the real-world data (i.e., Real Data) generated by SAMBO is well-aligned with the Offline Data and remains tightly concentrated, indicating the SAMBO’s effectiveness. In contrast, the Real Data from MOBILE deviates significantly from both the Offline and Model Data, still suffering from distribution shift, leading to diminished performance. We attribute this discrepancy to SAMBO’s robust handling of model bias and policy shift, ensuring that the Real Data, Model Data, and Offline Data remain closely aligned, effectively mitigating the distribution shift.
5.2 Benchmark Results
5.2.1 D4RL
Datasets. In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012). The evaluation encompasses 12 datasets spanning three environments (halfcheetah, hopper, walker2d) and four dataset types (random, medium, medium-replay, medium-expert). The “v2” versions of datasets are employed for assessment.
Baselines. We evaluate SAMBO against several baseline methods, including model-free methods such as BC (Bain & Sammut, 1995; Ross et al., 2011), which learns the behavior policy, CQL (Kumar et al., 2020), TD3+BC (Fujimoto & Gu, 2021), and EDAC (An et al., 2021), as well as model-based methods such as MOPO (Yu et al., 2020), COMBO (Yu et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023).
Table 2: Normalized average returns in NeoRL benchmark over 4 random seeds. The results for MOBILE are reported from the original paper, and MOBILE represents the results obtained by executing the official implementation.
Task Name BC CQL TD3+BC EDAC MOPO MOBILE MOBILE SAMBO (Ours)
HalfCheetah-L 29.1 38.2 30.0 31.3 40.1 54.7 54.8 39.8 $\pm$ 2.6
Hopper-L 15.1 16.0 15.8 18.3 6.2 17.4 18.2 19.4 $\pm$ 1.3
Walker2d-L 28.5 44.7 43.0 40.2 11.6 37.6 23.7 46.1 $\pm$ 5.7
HalfCheetah-M 49.0 54.6 52.3 54.9 62.3 77.8 79.0 62.5 $\pm$1.6
Hopper-M 51.3 64.5 70.3 44.9 1.0 51.1 43.4 67.6 $\pm$ 12.2
Walker2d-M 48.7 57.3 58.6 57.6 39.9 62.2 60.1 63.4 $\pm$ 4.0
HalfCheetah-H 71.3 77.4 75.3 81.4 65.9 83.0 71.8 71.6 $\pm$ 13.9
Hopper-H 43.1 76.6 75.3 52.5 11.5 87.8 42.3 93.9 $\pm$ 15.2
Walker2d-H 72.6 75.3 69.6 75.5 18.0 74.9 71.9 79.2 $\pm$ 0.6
(a) Hopper-L
(b) Hopper-M
(c) Hopper-H
Figure 4: Performance comparison of Hopper tasks in NeoRL benchmark.
Comparison Results. Results are presented in Table 1. Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al., 2020). SAMBO achieves strong performance on random datasets and demonstrates competitive performance across medium, medium-replay, and medium-expert datasets. This performance aligns with our theoretical analysis. By addressing both model bias and policy shift, SAMBO effectively balances the exploitation of offline data with the need to overcome the limitations of the behavioral policy. This approach enables the algorithm to be more aggressive and optimistic in its search for the optimal policy. In scenarios involving low-quality datasets, the aggressive exploration of SAMBO is more likely to yield benefits, potentially leading to the discovery of a superior policy. Conversely, a conservative approach may impede the learning of optimal policy in such contexts. For high-quality datasets, the advantages of aggressive exploration tend to diminish. Fortunately, in these scenarios, methods that emphasize estimating Q-value uncertainty perform effectively. This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
5.2.2 NEORL
Datasets. NeoRL (Qin et al., 2022) is a benchmark designed to emulate real-world scenarios by generating datasets through a more conservative policy, aligning closely with real-world data-collection scenarios. The datasets, characterized by their narrow scope and limited coverage, present a substantial challenge for offline RL algorithms. In this work, we examine nine datasets across three environments (HalfCheetah-v3, Hopper-v3, Walker2d-v3) and three dataset quality levels (L, M, H), corresponding to low, medium, and high-quality datasets. NeoRL offers different numbers of trajectories for each task (100, 1000, and 10000), and we consistently selected 1000 trajectories in our test.
Baselines. We compare SAMBO with baseline methods that align with the d4rl experiments in Section 5.2.1, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would require an excessive amount of computational time.
Comparison Results. Results are presented in Table 2 and Figure 4. We compare the learning curves of model-based approaches in the NeoRL benchmark. Since MOPO does not have an official implementation on NeoRL, we report its score based on the original NeoRL paper. The experiments show that Our algorithm achieves superior or competitive performance on the more challenging and realistic NeoRL datasets, which can be seen as lower-quality datasets compared to the D4RL datasets. In particular, SAMBO demonstrates exceptional performance in the Walker2d and Hopper environments, which, in contrast to the HalfCheetah environment, exhibit higher control complexity, more intricate action dimensions, and greater training difficulty. These offline datasets more accurately capture the dynamic control challenges encountered in real-world scenarios, making them significantly demanding and highlighting that SAMBO has strong potential for real-world applications. More experiments and discussions are provided in Appendix D.3.
(a) Ablation in halfcheetah-random dataset
(b) Tuning for $\alpha$
(c) Tuning for $\beta$
(d) Ablation study in Walker2d-H dataset
Figure 5: Illustration of hyperparameter experiments in D4RL benchmark and NeoRL benchmark. (a) and (d) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward. (b) A large model bias adjustment coefficient $\alpha$ increases conservativeness. (c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
5.3 Ablation Studies
@@
Ablation. Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification. To elucidate the importance of these components, we perform ablation studies on the Halfcheetah-random task in D4RL benchmark and Walker2d-H task in NeoRL benchmark. As shown in Figure. 5 (a) and (d), we assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent.
@@
Figure. 5 (a) show that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates. This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration. However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance. This negative effect is particularly evident in the narrow Walker2d-H dataset, as seen in Figure 5 (d). Conversely, the wo/ps ablation addresses model bias but neglects policy shift adjustments. While this reduces value misestimation, the algorithm remains constrained by the behavior policy due to the lack of policy shift modification. Consequently, wo/ps is less effective in discovering a more optimal policy. As shown in Figure 5 (a), the performance of wo/ps is inferior to sambo, and Figure 5 (d) illustrates that wo/ps converges slowly compared to sambo. In contrast, SAMBO incorporates not only model bias adjustment but also policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance.
Hyperparameter Tuning. SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$. The model bias adjustment coefficient $\alpha$ adjusts model-generated data, with larger values overly penalizing exploration and smaller values inducing over-optimism, both harming performance. To validate this, we fixed $\beta=0.01$ and varied $\alpha$ in the halfcheetah task. As shown in Fig. 5 (b), increasing $\alpha$ from 0 to 0.05 initially improves performance but later hinders it due to excessive conservativeness, highlighting the need for a balanced $\alpha$ in policy learning.
@@
The policy shift coefficient $\beta$ regulates modifications to offline data, controlling deviation from the behavior policy. Larger $\beta$ values enhance exploration by encouraging greater deviation from the behavior policy, but also increase the risk of unsafe or suboptimal actions. As shown in Fig. 5 (c), the algorithm exhibits performance fluctuations as $\beta$ increases from 0 to 0.1. Low $\beta$ may lead to suboptimal solutions, while high $\beta$ can induce excessive aggressiveness and instability. Therefore, $\beta$ also needs to be set appropriately.
6 Conclusion and Discussion
@@
In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift. Model bias results in distorted objectives, while policy shift obstructs the convergence of policies. We introduce a shifts-aware policy optimization framework to uniformly address these challenges through shift-aware rewards (SAR) and apply it within the model-based offline setting, thereby proposing our SAMBO-RL algorithm. Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms. Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training. In future work, We aim to explore alternative methods that directly address distribution shifts using datasets, thereby eliminating the need for classifier estimation. Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems. Our novel analysis and results related to policy shifts in this context open new avenues for better exploiting offline datasets and overcoming the implicit restrictions imposed by behavior policies.
Impact Statement
This study addresses the distribution shift problem in offline model-based reinforcement learning (RL), with the potential to facilitate the deployment of RL algorithms in real-world applications, particularly in scenarios where data collection is expensive. The research aims to enhance the robustness and reliability of RL systems in real world. However, the direct deployment of RL systems in high-risk domains, such as autonomous driving, may introduce substantial risks, especially if algorithms fail to adapt to unforeseen distribution shifts in real-world environments. Furthermore, the application of offline RL models, particularly in sensitive areas like healthcare, may raise significant ethical concerns. In such domains, the training data may not fully represent all patient groups. If distribution shift issues are not adequately addressed, the resulting treatment policies may fail for certain patient populations. Therefore, to ensure the safe, effective, and equitable deployment of RL systems in real world, rigorous testing, validation, and continuous monitoring are essential.
Appendix A Theoretical Analysis
A.1 The Supplement Analysis of Distribution Shift